OpenStudy (anonymous):

Simplify negative 4 and 1 over 4 − negative 9 and 1 over 2. negative 13 and 1 over 3 negative 5 and 1 over 4 13 and 1 over 3 5 and 1 over 4 HELP PLEASE AND IF YOU HELP ME I WILL GIVE YOU A METAL AND BECOME ONE OF YOUR FANS

OpenStudy (crashonce):

= 17/4 - (-19/2)

OpenStudy (crashonce):

= 17/4 + 19/2

OpenStudy (crashonce):

oh its -17/4sorry

OpenStudy (anonymous):

Can you tell me if its the 1st,2nd,3rd,4th @CrashOnce

OpenStudy (crashonce):

sorry but its D

OpenStudy (crashonce):

some calculation mistakes

OpenStudy (jack1):

\(\Large ~~~-4 \frac 14 - (-9 \frac 12)\) \(\Large = -4 \frac 14 - (-9 \frac 24)\) \(\Large = -( \frac {16}4+ \frac 14) - (-9 \frac 24)\) \(\Large = -( \frac {16}4+ \frac 14) - (- (\frac {36}4 + \frac 24))\) \(\Large = - \frac {17}4 + \frac {38}4\) \(\Large = - \frac {17}4 + \frac {38}4\) \(\Large = \frac {21}4\) \(\Large = \frac {20}4 + \frac {1}4\) \(\Large = 5 + \frac {1}4\)